OpenStudy (anonymous):

Convert 30 square ft to square meters, rounding to the nearest hundredths.

OpenStudy (anonymous):

Using 1 sq ft =0.092903 1/0.092903=x/30 x=30*1/0.092903 x=2.78709 x=2.79 sq meters

OpenStudy (anonymous):

Thank you.

OpenStudy (anonymous):

your welcome